calculate steady state temperature of the black body copper disc of radius 4cm and mass 10g. at surrounding temperature of 50 C . when kept initially at 25 C.Also determine stefan boltzmann constant
1
Expert's answer
2020-05-21T13:34:00-0400

The disk gains heat "Q_{in}" due to thermal conductivity from the surrounding and looses heat "Q_{out}" due to radiation.

Thus, according to the Stefan–Boltzmann law, the looses rate:


"\dfrac{dQ_{out}}{dt} = \pi r^2\sigma(T_{sur}^4 - T^4),"

where "T" is the current temperature of the disk, "\pi r^2" is the area of the surface of the disk, "\sigma = 5.67\cdot 10^{-8} \dfrac{W}{m^2K^4}" is the Stefan–Boltzmann constant and "T_{sur}" is the temperature of the surrounding in Kelvin.

The gain rate will be:


"\dfrac{dQ_{in}}{dt} = \dfrac{d}{dt}(mc(T - T_{in})) = mc\dfrac{dT}{dt},"

where "m" is the mass of the disk, "c" is the specific heat capacity of the cooper, "T_{in}" is the initial temperature of the disk in Kelvin.

In the steady state the rate of gain will be equal to the rate of loss, thus:


"\dfrac{dQ_{out}}{dt} = \dfrac{dQ_{in}}{dt}""\dfrac{dT}{dt} = \dfrac{\pi r^2\sigma}{mc}T_{sur}^4 - \dfrac{\pi r^2\sigma}{mc}T^4"


"\dfrac{dT}{dt} +\dfrac{\pi r^2\sigma}{mc}T^4 = \dfrac{\pi r^2\sigma}{mc}T_{sur}^4"

The solution of this equation is:


"T(t) = \dfrac{1}{\sqrt[3]{3\dfrac{\pi r^2\sigma}{mc}t + \dfrac{1}{(T_{start - T_{sur}})^3}}} + T_{sur}"

In steady state (as "t" goes to infinity):


"T_{steady} = T_{sur} = 273+50 = 323K"

Answer. 323 K or 50 C.